(2x+5)/5 = (3x+1)/2 + (-x+7)/2 Solve equation and check solution. Thank you!
Answer by linht117(37)

(2x+5)/5 = (3x+1)/2 + (-x+7)/2 rearrange equation
10*(3x+1)/2 + 10*(-x+7)/2 = 10*(2x+5)/5 find LCD-10 and multiply it to each term
10x+40=4x+10 solve for x
x= -5
______________________________
plug -5 original equation
[(2)(-5)]+5/5= [(3)(-5)]+1/2 + [(-(-5)+(7)]/2
-10+5/5 = -15+1/2 + 5+7/2
-5/5 = -14/2 + 12/2
-1 = -7 + 6
-1 = -1